Circular No. 7420

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

(719) ALBERT = 2000 JW_{8} An asteroidal object with unusual motion, first detected on May 1 by

J. A. Larsen and reobserved on May 3 and 6 by R. S. McMillan and J. V. Scotti in the course of the Spacewatch program at Kitt Peak and placed on the NEO Confirmation Page, was also observed on May 9 by M. Hicks and R. Fevig with the 2.1-m Kitt Peak reflector. G. V. Williams, Minor Planet Center, has identified this object, designated 2000 JW_{8}, with (719) Albert, the last ‘lost’ numbered minor planet. Selected observations follow:

2000 UT α_{2000} δ_{2000} m_{1} Observer
May 1.19935 11^{h}38^{m}04.^{s}81 +2^{o}35′21′′.7 21.6 Larsen
2.17479 11 37 40.78 +2 39 31.5 21.1 McMillan
6.20445 11 36 12.18 +2 55 25.2 22.3 Scotti
9.16403 11 35 18.54 +3 05 46.0  Hicks

Further details, including residuals for the 1911 and 2000 observations, are given on MPEC 2000-J37, and Williams’ orbital elements there, reproduced below, show that the analysis by Kristensen and West (1989, A.Ap. 218, 317; a = 2.645 AU, e = 0.550, i = 11^{o}.0) was a significant improvement over earlier estimates. The 1913 Crossley search (Curtis 1913, Lick Obs. Bull. 8, 47) was centered ∼ 2^{o} from the object’s actual position, and the 1915 prediction (Seagrave 1915, A.J. 28, 202) requires correction by ΔT = +47 days. Close (September) approaches to the earth occur in 1911, 1941, 1971 and 2001 (minimum 0.205, 0.206, 0.316 and 0.286 AU, and, with H = 15.8, brightest V = 13.7, 13.7, 15.5 and 15.2, respectively).

Epoch = 1911 Dec. 30.0 TT

T = 1911 Aug. 29.5289 TT  ω = 151.^{o}9113
e = 0.548973  Ω = 186.8071 }2000.0
q = 1.189954 AU  i = 10.9429

a = 2.638322 AU n^{o} = 0.2299917 P = 4.285 years Epoch = 2000 Feb. 26.0 TT

T = 2001 Aug. 21.2582 TT  ω = 154.^{o}2885
e = 0.548273  Ω = 184.9305 }2000.0
q = 1.191053 AU  i = 11.3095
a = 2.636667 AU n^{o} = 0.2302084 P = 4.281 years

There were also perihelia T = 1941 Aug. 31.0 and 1971 Aug. 20.0.

2000 May 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Brian G. Marsden